a rental car place rents out cars. barney is charged $111 for 3 days and 300 miles, mary is charged $207 for 5 days and 600 miles. what do they charge per day and per mile?
Answer by checkley77(12844)

3D+300M=111 MULTIPLY BY 2
5D+600M=207 MULTIPLY BY -1 & ADD
6D+600M=222
-5D-600M=-207
----------------
D=$13 PER DAY
3*15+300M=111
45+300M=111
300M=111-45
300M=66
M=66/300
M=$.22 PER MILE.
PROOF:
5*15+600*.22=207
75+132=207
207=207